OpenStudy (anonymous):

please help me solve this problem x^2 -3x +________identify the value that would make the following a perfect square

OpenStudy (anonymous):

2.25 That would make it (x-1.5)^2.

OpenStudy (anonymous):

Thank you for your help

OpenStudy (anonymous):

The idea is, you want to have \((x-a)^2\). Expanded, you have \(x^2-2ax+a^2\). Matching up with the given polynomial: \[x^2-3x+\cdots=x^2-2ax+a^2\] So you have \[\begin{cases}-3=-2a\\ \cdots=a^2\end{cases}\] You'd solve the first equation for \(a\), then plug it into the second equation.

OpenStudy (anonymous):

okay so I need to work this problem out

OpenStudy (anonymous):

If you're told to show work, then this is the way to do it.

OpenStudy (anonymous):

okay thank you

OpenStudy (anonymous):

You're welcome

OpenStudy (anonymous):

one more questions my class mate came up with 9/4 is that correct

OpenStudy (anonymous):

how do I work the second part of this problem

OpenStudy (anonymous):

9/4 = 2.25